OpenStudy (yacoub1993):

Li went for a mountain-bike ride in a relatively flat wooded area. She rode for 6 km in one direction and then turned and pedaled 16 km in another. Finally she turned in the direction of her starting point and rode 8 km. When she stopped, was it possible that Li was back at her starting point? Explain.

OpenStudy (potatoes.ramu):

No it's not possible! Using pythagoras the theorem! adding the squares of any of the two numbers doesn't give you the square of the other number! :)

OpenStudy (yacoub1993):

@Potatoes.ramu so what should i write, i mean how do i explain

OpenStudy (potatoes.ramu):

Say it's not possible. And then squares of two numbers do not equal the square of the third one! like 6^2+8^2=36+64=100 is not equal to 16^2